Find the least 5- digit number which is exactly divisible by 20 , 25 , 30.

Answer by Theo(13342)

the least common multiple of 20, 25, and 30, is 300.
i believe that any larger numbers that are multiples 20, 25, and 30 have to be multiples of 300.
10,000 / 300 = 33.333
33 * 300 = 9900
34 * 300 = 10200
that looks like the smallest 5 digit number that's divisible by 20, 25, 30 is 10200.
i tested this out with a program that i wrote in excel and it tells me that my assumption is correct and that 10200 is the least 5 digit number that is a common multiple of 20, 25, and 30.
the least 5 digit number has to be a number greater than or equal to 10000.
the next smaller number is 9999 which is a 4 digit number.